ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 83, 1993

Bollobás, B., C. Gotsman and E. Shamir Connectivity and dynamics for random subgraphs of the directed cube

321

Bose, C. Mixing examples in the class of piecewise monotone and continuous maps of the unit interval

129

Brendle, J. and H. Judah Perfect sets of random reals  153

Carter, J. E. See S. A. Vanstone

Carter, M. W. See S. A. Vanstone

Colbourn, C. J. See S. A. Vanstone

Donder, H.-D. and P. Matet Two cardinal versions of diamond  1

Garcia-Cuerva, J., R. Macias and J. L. Torrea The Hardy– Littlewood property of Banach lattices

177

Gelbart, S., J. Rogawski and D. Soudry On periods of cusp forms and algebraic cycles for U(3)

213

Gotsman, C. See B. Bollobás

Grzeszczuk, P. and J. Matczuk Goldie conditions for constants of algebraic derivations of semiprime algebras

329

Iwanik, A., M. Lemańczyk and D. Rudolph Absolutely continuous cocycles over irrational rotations

73

Jacob, B. and A. R. Wadsworth Division algebras with no common subfields

353

Judah, H. See J. Brendle

Ledger, A. J. Congruence theorems for Riemannian submanifolds  295

Lemańczyk, M. See A. Iwanik

Macias, R. See J. Garcia-Cuerva

Matczuk, J. See P. Grzeszczuk

Matet, P. See H.-D. Donder

Montgomery, S. Biinvertible actions of Hopf algebras  45
Mountford, T. S. A solution to a problem of Dubins and Savage  253
Perets, G. S. d-Critical modules of length 2 over Weyl algebras  361
Rao, M. and H. Šikić Potential inequality  97

Rees, R. See S. A. Vanstone Rogawski, J. See S. Gelbart Rosa, A. See S. A. Vanstone Rudolph, D. See A. Iwanik Schellenberg, P. J. See S. A. Vanstone Shamir, E. See B. Bollobás

Shelah, S. Number of open sets for a topology with a countable basis 369

Šikić, H. See M. Rao Soudry, D. See S. Gelbart

Steprāns, J. Martin’s Axiom and the transitivity of P_{c}-points  257

Stinson, D. R. See S. A. Vanstone Torrea, J. L. See J. Garcia-Cuerva S. A. Vanstone, D. R. Stinson, P. J. Schellenberg, A. Rosa,

R. Rees, C. J. Colbourn, M. W. Carter and J. E. Carter Hanani triple systems

305

Wadsworth, A. R. See B. Jacob

Walther, H. Longest cycles in polyhedral graphs  203
Weston, A. Some non-uniformly homeomorphic spaces  375

Xiao Pingan See Zhou Ji Zakrzewski, P. The existence of invariant σ-finite measures for a

group of transformations

275

Zakrzewski, P. The existence of invariant probability measures for a group of transformations

343

Zhou Ji and Xiao Pingan A note on the valence of certain means  289